im a parent trying to help my son figure out the process of getting to the right answer for this equation, 9b+6-5b=18 then the teacher shows b=3 i did the problem just like the example but can't figure out where the 3 comes from i was able to see how the answer for the example came about. Another equation is 2y-7+y=8 please help so i can explain to my son thanks
Answer by nerdybill(7384)

im a parent trying to help my son figure out the process of getting to the right answer for this equation, 9b+6-5b=18 then the teacher shows b=3 i did the problem just like the example but can't figure out where the 3 comes from i was able to see how the answer for the example came about. Another equation is 2y-7+y=8 please help so i can explain to my son thanks
.
9b+6-5b=18
combining b's on the left:
4b+6=18
subtracting 6 from both sides:
4b+6-6=18-6
4b=12
dividing both sides by 4:
(4b)/4 = 12/4
b = 3
.
2y-7+y=8
combining y's on the left:
3y-7=8
add 7 to both sides:
3y-7+7=8+7
3y=15
dividing both sides by 3:
(3y)/3=15/3
y = 5